OpenStudy (idealist10):

Solve the initial value problem (x^2-4)y"+4xy'+2y=x+2, y(0)=-1/3, y'(0)=-1, given that y1=1/(x-2) satisfies the complementary equation.

OpenStudy (idealist10):

y=u(x-2)^-1 y'=-u(x-2)^(-2)+u'(x-2)^-1 y"=2u(x-2)^(-3)-u'(x-2)^(-2)-u'(x-2)^(-2)+u"(x-2)^-1 Subbing: \[(x^2-4)y''+4xy'+2y=u''(x+2)-\frac{ 2u'(x+2) }{ x-2 }\]...

OpenStudy (idealist10):

Simplify, \[u''+\frac{ 2 }{ x+2 }u'=1\] z=u' subbing: \[z'+\frac{ 2 }{ x+2 }z=1\] \[(x+2)^2z=\frac{ (x+3)^3 }{ 3 }+C\] \[z=u'=\frac{ x+2 }{ 3 }+C(x+2) ^{-2}\] integrate \[u=\frac{ x^2 }{ 6 }+\frac{ 2x }{ 3 }-C _{1}(x+2) ^{-1}+C _{2}\] Since \[y=\frac{ u }{ x-2 }\], \[y=\frac{ x^2+4x }{ 6(x-2) }-\frac{ C _{1} }{ x^2-4 }+\frac{ C _{2} }{ x-2 }\] plug in y(0)=-1/3, -1/3=C1/4-C2/2 At the end, I didn't get the right answer. The answer in the book says \[y=\frac{ (x+2)^2 }{ 6(x-2) }+\frac{ 2x }{ x^2-4 }\]

OpenStudy (idealist10):

@Kainui @Hero

OpenStudy (kainui):

Are you sure they're really different and not just different ways of arranging them by algebra?

OpenStudy (idealist10):

Well, the constants I got are C1=4 and C2=8/3. I don't know if those are right and if the book's answer is right.

OpenStudy (idealist10):

@freckles

OpenStudy (freckles):

i have to go with @kainui on this one... have you checked to see if both answers are the same... try to write their answer as your answer \[y=\frac{(x+2)^2}{6(x-2)}+\frac{2x}{x^2-4} \\ y=\frac{x^2+4x+4}{6(x-2)}+\frac{2x}{x^2-4} \\ y=\frac{x^2+4x}{6(x-2)}+ \frac{4}{6(x-2)}+\frac{2x}{x^2-4} \\ y=\frac{x^2+4x}{6(x-2)}+\frac{2}{3(x-2)}+\frac{1}{x-2}+\frac{1}{x+2} \\ y=\frac{x^2+4x}{6(x-2)}+\frac{2}{3(x-2)}+\frac{6}{3(x-2)} -\frac{1}{x-2}+\frac{1}{x+2} \\ y=\frac{x^2+4x}{6(x-2)} +\frac{8}{3(x-2)}+\frac{1}{x+2}-\frac{1}{x-2}\] ... see if you can continue there is only one more step really

OpenStudy (freckles):

though you could have integrated differently to get their answer without undoing and doing stuff to see if their answer is your answer

OpenStudy (freckles):

\[u'=\frac{x+2}{3}+C(x+2)^{-2} \\ u=\frac{(x+2)^2}{3 \cdot 2}-C(x+2)^{-1} +D \\ u=\frac{(x+2)^2}{6}-C(x+2)^{-1}+D \\ \text{ instead of writing } \\ u'=\frac{x}{3}+\frac{2}{3}+C(x+2)^{-2} \\ \text{ and the integrating which gives } \\ u=\frac{x^2}{6}+\frac{2}{3}x-C(x+2)^{-1}+D\] neither way is wrong

OpenStudy (idealist10):

But are the constants I got right or wrong? I got C2=8/3 and C1=4, when I plug them into y, I got a weird answer.

OpenStudy (freckles):

you don't know where to continue from the book's answer above?

OpenStudy (freckles):

\[y=\frac{(x+2)^2}{6(x-2)}+\frac{2x}{x^2-4} \\ y=\frac{x^2+4x+4}{6(x-2)}+\frac{2x}{x^2-4} \\ y=\frac{x^2+4x}{6(x-2)}+ \frac{4}{6(x-2)}+\frac{2x}{x^2-4} \\ y=\frac{x^2+4x}{6(x-2)}+\frac{2}{3(x-2)}+\frac{1}{x-2}+\frac{1}{x+2} \\ y=\frac{x^2+4x}{6(x-2)}+\frac{2}{3(x-2)}+\frac{6}{3(x-2)} -\frac{1}{x-2}+\frac{1}{x+2} \\ y=\frac{x^2+4x}{6(x-2)} +\frac{8}{3(x-2)}+\frac{1}{x+2}-\frac{1}{x-2} \] hint try combing the last two fractions

OpenStudy (idealist10):

It's 2x/(x^2-4).

OpenStudy (freckles):

combining the last two fractions?

OpenStudy (freckles):

\[\frac{1}{x+2}-\frac{1}{x-2}=\frac{(x-2)-(x+2)}{(x+2)(x-2)}=\frac{x-x-2-2}{x^2-4}=\frac{-4}{x^2-4}\]

OpenStudy (freckles):

you should see the book's answer is not equivalent to your answer

OpenStudy (freckles):

\[y=\frac{x^2+4x}{6(x-2)}+\frac{8}{3(x-2)}-\frac{4}{x^2-4} \\ =\frac{(x+2)^2}{6(x-2)}+\frac{2x}{x^2-4}\]

OpenStudy (freckles):

as shown above

OpenStudy (idealist10):

You know what? I got the right answer. I found my mistake. \[u'=\frac{ x+2 }{ 3 }+C(x+2) ^{-2}\] integrate, it should be \[u=\frac{ (x+2)^2 }{ 6 }-C _{1}(x+2) ^{-1}+C _{2}\] , not \[u=\frac{ x^2 }{ 6 }+\frac{ 2x }{ 3 }-C _{1}(x+2) ^{-1}+C _{2}\]

OpenStudy (freckles):

you didn't make a mistake

OpenStudy (freckles):

I showed their answer is same as yours above :p

OpenStudy (freckles):

also you can write (x+2)/3 as x/3+2/3 and integrate if you want or straight up integrate (x+2)/3

OpenStudy (idealist10):

I just want to say, thank you! :)

OpenStudy (freckles):

np

OpenStudy (freckles):

i see why there was confusion now "you should see the book's answer is not equivalent to your answer " this sentence was suppose to say now equivalent instead of not equivalent but anyways the work alone should have reflected the now instead of the not :p